We had good beam all night with currents up to 99mA, and the ABS spin was 
flipped every 10 seconds.  The nozzle and cell temperatures were stable 
all night (70K and 96K).  Ligit was constant at 1.25x 10^-7 Torr.
>From the previous shift (or earlier?) the HV GUI was left in the state of 
injection permitted, which allows CCR to inject manually with the high 
voltages on.  The Injection (BLAST) override button should normally be 
"off".  Unfortunately we had two fills with the high voltages on, which 
was after a few fills where the voltages were ramped.  CCR had been 
injecting automatically until this point.  We turned the WC's off for 
25min after the injection flash, and the signal to noise was the same in 
runs before and after.
Tancredi asked for a summary of good runs over the past day, so including 
this shift, we have the run numbers:
723-727, 730 (Sat B)
732-733 (Free of beam gate problem, Sat B)
734-735, 737-739, 741-744, 746-747, 751, 753-759, 761 (Sat C & Sun A)
We also have another 4 runs (Sun A) that caused segmentation faults in 
lrn.C.  The runs from 746 on were analyzed using "root charge.C -ltmax 
0.995 -ltmin".
YX and BC.
This archive was generated by hypermail 2.1.2 : Mon Feb 24 2014 - 14:07:29 EST